# Agent-X: Evaluating Deep Multimodal Reasoning in Vision-Centric Agentic Tasks

### v1

@@ 3.3 Human-Guided Refinement
Raw queries / Refined queries 1021 / 828
@@
Figure 3: Overview of the Agent-X benchmark. (a) Key data statistics. (b) Overall frequency of tool usage and number of steps. (c) Distribution of tasks across the six vision-centric environments.
Task Expansion and Validation. After establishing the initial queries, we adopt a diversified expansion approach to scale the tasks. Annotators are instructed to edit generated queries by varying the scenarios and maintaining or modifying each query’s toolset $\mathcal{T}$. This approach promotes query diversity without deviating from the core functionality of the tools. Detailed instructions for query refinement are provided in Appendix §G. Annotators are encouraged to reimagine the scenarios creatively, ensuring a wide range of query types and complexity levels. Following the query expansion, a rigorous quality assurance process is conducted. The second iteration reviews the modified queries to ensure they require a multi-step deep reasoning trace and maintain logical consistency. Queries that fail to meet the quality criteria are filtered out. For web search queries, we implement two safeguards: (1) The query must be answerable using real-time search results rather than relying solely on static knowledge from the language model, and (2) it should reference credible information sources when necessary, such as “What were the global average temperatures reported in April 2024?” rather than general queries like “What is climate change?”.
Tool Chain Construction and Validation. After refining the queries, we provide each query $\mathcal{Q}$ and its corresponding visual input $\mathcal{V}$ to an LMM, along with the toolset $\mathcal{T}_{c}$, which generates an initial reasoning trace, final answer $\mathcal{A}$, and a justification $\mathcal{J}$. The LMM’s output specifies the sequence of tool invocations, the input arguments for each tool, and the resulting outputs, formatted in a JSON-style dialogue sequence. Human annotators then rigorously review and refine the LMM-generated trace after this automated generation to ensure that the selected tools, input arguments, and intermediate reasoning steps are logically consistent and accurate. They correct factual errors, replace inappropriate tools, and verify that the final answer $\mathcal{A}$ aligns with the reasoning trace. We exclude the queries that lack logical consistency or cannot be solved reliably to maintain overall quality and reliability. Appendix §K illustrates examples of LMM-generated queries and reasoning traces before and after human refinement.
@@ 3.4 Dataset Composition and Statistics
The final Agent-X benchmark comprises a diverse set of queries covering multiple categories, including factual, interpretive, and generative tasks. The dataset is constructed using 14 executable tools, spanning six diverse environments, $e.g.$, autonomous driving[15, 64, 74], security and surveillance[4, 53, 35, 42], math reasoning[36, 31, 14], web navigation[54, 80], sports[25, 65], and generic visual reasoning [30, 62]. These categories are designed to support a wide range of multimodal reasoning tasks. The total number of tasks, query analysis, tool analysis, and annotation labor hours are detailed in Figure 3. Most queries involve the use of two to four executable tools. This diverse setup ensures that Agent-X can comprehensively evaluate model performance across various tool-augmented reasoning tasks. A detailed Datacard describing the dataset’s purpose, structure, and collection process is included in Appendix §A.
@@ 4.1 Experimental Setup
Table 3: Evaluation Metrics categorized by Mode. This table outlines the full suite of metrics used in Agent-X benchmark, organized by Step-by-Step, Deep Reasoning, and Outcome modes.
Metric Notation Description
Step-by-Step Mode
Grounding Score $\mathbf{G_{s}}$ Correct reference to objects, regions, or attributes in the input.
Tool Precision $\mathbf{T_{p}}$ Accuracy of selecting the correct tool at each reasoning step.
Tool Accuracy $\mathbf{T_{a}}$ Correct use of tools with appropriate inputs and outputs.
Deep Reasoning Mode
Faithfulness $\mathbf{F_{acc}}$ Logical consistency across the reasoning process.
Context Score $\mathbf{C_{s}}$ Effective use of multimodal and commonsense context.
Factual Precision $\mathbf{F_{p}}$ Correctness of factual information without hallucination.
Semantic Accuracy $\mathbf{S_{acc}}$ Coverage of all semantically necessary elements.
Outcome Mode
Goal Accuracy $\mathbf{G_{acc}}$ Final answer accuracy for factual and interpretive queries.
Goal Accuracy w/ImgGen $\mathbf{G^{*}_{acc}}$ Final answer accuracy for generative queries.
Toolset Accuracy $\mathbf{T_{acc}}$ F1 score for overall correct tool selection and use.
Metrics: We design a suite of fine-grained metrics (Table 3) to evaluate agentic reasoning across all three modes comprehensively. These metrics capture key aspects of the agent pipeline. For Goal Accuracy, we exclude image generation queries and focus on factual and interpretive queries, using exact matching against the gold answer and descriptive matching against labeled responses using GPT-4o [21] respectively. For image generation, we define Goal Accuracy/ImgGen, which assesses the correctness of predicted input parameters, assuming accurate inputs yield suitable images. Detailed metrics computation and implementation are provided in Appendix §C.
@@ 4.2 Benchmark Results
Table 4 presents the core evaluation results of Agent-X across three modes: Step-by-Step, Deep Reasoning, and Outcome. We evaluate a mix of closed-source models (e.g., GPT [2, 45] and Gemini [55, 56]) and strong open-source counterparts (e.g., InternVL [12, 81] and Qwen [6, 8]). While closed-source models lead overall, open-source models show competitive behavior in select metrics, offering key insights across agentic frameworks, which we summarize below. We additionally report results using two alternative evaluation judges, Qwen and human annotators, in Appendix §E. The evaluation prompts provided to GPT-4o, Qwen, and human annotators are included in Appendix §J.
Insight 1: Real-world tool-use tasks remain challenging for current LMM agents. Despite multimodal capabilities, no model exceeds 50% in Goal Accuracy ($\mathbf{G_{acc}}$). o4-mini, the best performer, achieves only 45%, while most open-source models are below 30% highlighting the difficulty of grounding, tool use, and final answer consistency in real-world settings.
@@
Phi-4-VL-Instruct [1] 0.13 0.21 0.24 0.61 0.19 0.47 0.40 0.11 0.26 0.42
InternVL2.5-8B [12] 0.45 0.31 0.47 0.68 0.47 0.52 0.60 0.28 0.55 0.58
Gemma-3-4B [57] 0.26 0.30 0.78 0.61 0.54 0.38 0.54 0.27 0.67 0.60
InternVL3-8B [81] 0.46 0.34 0.54 0.68 0.45 0.70 0.40 0.20 0.59 0.62
VideoLLaMA3-7B [75] 0.45 0.28 0.46 0.65 0.46 0.62 0.54 0.28 0.54 0.54
Qwen2.5-VL-7B [6] 0.54 0.43 0.63 0.75 0.57 0.56 0.67 0.36 0.65 0.67
@@
Gemini-1.5-Pro [56] 0.43 0.23 0.84 0.62 0.45 0.53 0.62 0.04 0.56 0.48
Gemini-2.5-Pro [55] 0.40 0.36 0.81 0.72 0.48 0.64 0.73 0.40 0.56 0.62
GPT-4o [21] 0.60 0.47 0.72 0.81 0.57 0.79 0.59 0.37 0.70 0.68
OpenAI-o4-mini [45] 0.42 0.32 0.89 0.71 0.51 0.60 0.80 0.45 0.67 0.63
*Results for the additional models are presented in Appendix §D.
Insight 2: Strong reasoning capabilities contribute to higher task success rates. Models that maintain consistently strong scores across reasoning metrics are likelier to perform well on final task outcomes. GPT-4o, for instance, achieves high scores in $\mathbf{F_{acc}}=0.81$, $\mathbf{F_{p}}=0.79$, and $\mathbf{S_{acc}}=0.59$, which correlates with a relatively high $\mathbf{G_{acc}}=0.37$ and $\mathbf{T^{s}_{acc}}=0.68$. Likewise, Qwen2.5, with solid reasoning scores ($\mathbf{C_{s}}=0.57$, $\mathbf{S_{acc}}=0.67$), attains $\mathbf{G_{acc}}=0.36$, outperforming many open-source peers. These results indicate that deep reasoning and structured execution significantly enhance final task success in vision-centric agentic tasks.
@@ 4.3 Error Analysis
Figure 4: Qualitative comparison of GPT-4o and VideoLLaMA3-7B on visual reasoning tasks from Agent-X. GPT-4o hallucinates tool use and produces wrong justifications, while VideoLLaMA3-7B lacks temporal reasoning and fails to align observations across frames. Both models struggle with tool grounding and structured reasoning. Comparisons of other models are provided in Appendix §F.2.
Qualitative Analysis: We conduct a qualitative error analysis (Figure 4) to uncover key limitations in current vision-centric agentic models. Our findings reveal four prevalent issues: (1) agents often exhibit shallow reasoning by skipping frame-wise or multi-step analysis, especially in video-based tasks; (2) models frequently misuse or hallucinate tools not defined in the metadata, indicating poor tool grounding; (3) many responses violate expected output formats, producing incomplete or non-JSON-compliant traces that hinder evaluation; and (4) agents tend to hallucinate reasoning steps or bypass visual verification, resulting in factual inaccuracies. These insights highlight the need for better temporal reasoning, tool schema adherence, structured output enforcement, and visually grounded planning in future agentic systems.
Table 5: Error Breakdown Across Models. Summary of common planning, formatting, and reasoning errors on Agent-X benchmark across GPT-4o, Gemini-1.5-Pro, and InternVL3-8B. Extended error breakdowns for additional models are included in Appendix §F.1.
Error Type GPT-4o [21] Gemini-1.5-Pro [55] InternVL3-8B [81]
@@
Quantitative Analysis: We analyze agentic failure modes by categorizing key error types across three representative models: GPT-4o [21], Gemini-1.5-Pro [55], and InternVL3-8B [81]. Table 10 highlights notable behavioral contrasts. GPT-4o is generally structure-aware, with relatively few format-related errors (e.g., 13.2% multiple tool calls, 6.7% final formatting issues). However, it often fails to act. 17.6% of its errors involve no response, and another 18.5% result from incorrect visual interpretation implying that GPT-4o plays safe but hesitates to act. In contrast, Gemini-1.5-Pro exhibits a different profile; it is highly error-prone in format compliance, with 44.5% of its errors due to malformed JSON arguments, the highest among all models, and 10.3% from final answer formatting issues. Notably, it also struggles with visual understanding (34.3% misinterpretations) but rarely avoids action (just 0.2% non-responses), indicating Gemini-1.5-Pro is aggressive but format-fragile. InternVL3-8B presents a more balanced failure mode, though still marked by high format issues (33.8% JSON errors, 16.4% final formatting). Like GPT-4o, it struggles with spatial reasoning (13.5%) and visual misinterpretation (14.1%), suggesting InternVL3-8B falls in between but suffers from both structural and perceptual challenges.
Figure 5: Tool Call analysis across LMM families on Agent-X benchmark.
Tool Call Success and Failure Analysis. Figure 5 shows the number of successful and failed tool invocations per model. A successful call executes without format or argument errors. The GPT family achieves the highest success rate (83.8%) relative to its total calls, with GPT-4o demonstrating a strong balance between usage and reliability. InternVL models lead to absolutely successful calls but exhibit high failure counts, primarily due to invalid tool usage. Qwen2.5-VL-8B shows the highest precision, with only 109 failures out of 2241 calls, reflecting robust format compliance. In contrast, o4-mini is the most aggressive, issuing 3374 calls but suffering the highest failure rate (1038). Gemini-3-4B and VideoLLaMA3-7B show moderate tool use, with Gemini-3-4B maintaining a lower failure count. These trends highlight three behavioral patterns: aggressive models ($e.g.$, InternVL3-8B) invoke tools frequently but are prone to errors; conservative models ($e.g.$, Qwen2.5-VL-7B) prioritize accuracy over volume; and balanced models ($e.g.$., GPT-4o) combine moderate tool use with high reliability. For additional analyses on how agent performance varies by task difficulty and reasoning depth, please refer to Appendix §F.3.
@@ 5 Conclusion
We present Agent-X, a comprehensive benchmark for evaluating the reasoning capabilities of vision-centric agents. The tasks in Agent-X are context-rich and span diverse multimodal scenarios. Our evaluation framework includes executable tools categorized across six agentic environments, ensuring a wide range of task coverage. We evaluate 10 state-of-the-art LMMs on Agent-X and find that even strong closed-source models struggle with agentic tasks. Our insights provide actionable guidance for improving agentic capabilities. We believe that Agent-X will drive future research in enhancing multimodal reasoning and robust tool integration for vision-centric agents.
Limitations. Agent-X is currently monolingual, limiting its use for evaluating multilingual agents. While it spans six diverse environments, potential distributional or cultural biases may still affect fairness and generalization. Additionally, the semi-automated query and tool-chain generation process, though cost-efficient and requiring less human effort, can occasionally produce lower-quality samples compared to fully manually annotated data.
@@
Section A Agent-X Data Card
@@
Section D Evaluation of newly–added models
@@
Section G Instruction for Annotators
Section H Agent-X Task Examples
Section I Generation Prompt Details
Section J Evaluation Prompt Details
Section K Benchmark Curation Examples

### v2

@@ 3.3 Human-Guided Refinement
Raw tasks / Refined tasks 1021 / 828
@@
Figure 3: Overview of the Agent-X benchmark. (a) Key data statistics. (b) Overall frequency of tool usage and number of steps. (c) Distribution of tasks across six environments.
Task and Reasoning Construction. Our task construction process begins with an automated generation phase, where an LMM produces three candidate queries for each of 1,021 initial visual inputs (images or video frames). Annotators then select the best query from each set, yielding 1,021 raw queries. Detailed annotation guidelines, including query construction rules and reasoning/answer refinement protocols, are provided in Appendix §H. Each candidate is carefully reviewed for clarity, realism, and suitability to ensure that it cannot be answered directly from the input alone, but instead necessitates meaningful tool use, is expandable to diverse scenarios, and supports multi-step reasoning. Through this refinement process, queries failing to meet these criteria are discarded, resulting in a final pool of 828 validated tasks. Each task is paired with one or more unique visual inputs, with no visual input reused across tasks. For web search queries, we additionally enforce two safeguards: (1) the query must be answerable using real-time search rather than relying solely on static knowledge, and (2) it should reference credible sources, such as “What were the global average temperatures reported in April 2024?”, rather than overly general prompts like “What is climate change?”
In the second stage, we construct the toolchain for each validated task. Each query $\mathcal{Q}$ and its corresponding visual input $\mathcal{V}$ are provided to an LMM along with the toolset $\mathcal{T}_{c}$, which generates an initial reasoning trace, final answer $\mathcal{A}$, and justification $\mathcal{J}$. The reasoning trace specifies the sequence of tool invocations, input arguments, and intermediate outputs, formatted in a JSON-style dialogue. Human annotators rigorously review and refine these LMM-generated traces to ensure logical consistency, correctness of tool use, and factual alignment of the final answer $\mathcal{A}$ with the reasoning process. Annotators correct errors, replace inappropriate tool choices, and filter out tasks that cannot be solved reliably. Appendix §L illustrates examples of LMM-generated queries and reasoning traces before and after human refinement.
@@ 3.4 Dataset Composition and Statistics
The final Agent-X benchmark comprises a diverse set of queries covering multiple categories, including factual, interpretive, and generative tasks. The dataset is constructed using 14 executable tools, spanning six diverse environments, $e.g.$, autonomous driving (Cordts et al., 2016; Wood, 2020; Yu et al., 2020), security and surveillance (Aktı et al., 2019; Sultani et al., 2018; Lu et al., 2013; Naphade et al., 2023), math reasoning (Lu et al., 2023; Ling et al., 2017; Cobbe et al., 2021), web navigation (Sunkara et al., 2022; Zhou et al., 2023), sports (Karpathy et al., 2014; Wu et al., 2022), and generic visual reasoning (Lin et al., 2014; Wang et al., 2024). Full details of the source datasets are provided in Appendix §C.1.
The annotation effort amounted to roughly 50 hours per annotator, reflecting the full human refinement and validation process. Approximately 800K API tokens were used during benchmark construction. These categories are designed to support a wide range of multimodal reasoning tasks. The total number of tasks, query analysis, tool analysis, and annotation labor hours are detailed in Figure 3. Most queries involve the use of two to four executable tools. This diverse setup ensures that Agent-X can comprehensively evaluate model performance across various tool-augmented reasoning tasks.
@@ 4.1 Experimental Setup
Table 3: Evaluation Metrics. This table outlines the full suite of metrics used in Agent-X benchmark, organized by Step-by-Step, Deep Reasoning, and Outcome modes.
Metric (Symbol) Description
Step-by-Step Mode Grounding Score ( $\mathbf{G_{s}}$ ) Correct reference to objects, regions, or attributes in the input.
Tool Precision ( $\mathbf{T_{p}}$ ) Accuracy of selecting the correct tool at each reasoning step.
Tool Accuracy ( $\mathbf{T_{a}}$ ) Correct use of tools with appropriate inputs and outputs.
Deep      Reasoning Mode Faithfulness ( $\mathbf{F_{acc}}$ ) Logical consistency across the reasoning process.
Context Score ( $\mathbf{C_{s}}$ ) Effective use of multimodal and commonsense context.
Factual Precision ( $\mathbf{F_{p}}$ ) Correctness of factual information without hallucination.
Semantic Accuracy ( $\mathbf{S_{acc}}$ ) Coverage of all semantically necessary elements.
Outcome Mode Goal Accuracy ( $\mathbf{G_{acc}}$ ) Final answer accuracy for factual and interpretive queries.
Goal Accuracy w/ImgGen ( $\mathbf{G^{*}_{acc}}$ ) Final answer accuracy for generative queries.
Toolset Accuracy ( $\mathbf{T_{acc}}$ ) F1 score for overall correct tool selection and use.
Metrics: We design a suite of fine-grained metrics (Table 3) to evaluate agentic reasoning across all three modes comprehensively. These metrics capture key aspects of the agent pipeline. For Goal Accuracy, we exclude image generation queries and focus on factual and interpretive queries, using exact matching against the gold answer and descriptive matching against labeled responses using GPT-4o (Hurst et al., 2024) respectively. For image generation, we define Goal Accuracy/ImgGen, which assesses the correctness of predicted input parameters, assuming accurate inputs yield suitable images. Detailed metrics computation and implementation are provided in Appendix §C.
@@ 4.2 Benchmark Results
Table 4 presents the core evaluation results of Agent-X across three modes: Step-by-Step, Deep Reasoning, and Outcome. We evaluate a mix of closed-source models (e.g., GPT (Achiam et al., 2023; OpenAI, 2025) and Gemini (Team et al., 2023; 2024a)) and strong open-source counterparts (e.g., InternVL (Chen et al., 2024a; Zhu et al., 2025) and Qwen (Bai et al., 2023b; c)). While closed-source models lead overall, open-source models show competitive behavior in select metrics, offering key insights across agentic frameworks, which we summarize below. Beyond these core comparisons, we further extend the benchmark with additional open-source vision LMMs, as detailed in Appendix §D. We additionally report results using two alternative evaluation judges, Qwen and human annotators, in Appendix §E. The evaluation prompts provided to GPT-4o, Qwen, and human annotators are included in Appendix §K. To mitigate bias, all predictions are cross-checked by multiple graders (GPT-4o, Qwen-14B, and humans), with consistent model rankings across settings. Human scores correlate strongly with automatic grading, and residual discrepancies are uniformly distributed, confirming no systematic favoritism. Furthermore, evaluation metrics are explicitly bias-aware (decoupling syntax from semantics, normalizing tool arguments), and task seeds were rewritten under strict QA to avoid leakage. Together, these safeguards ensure fairness, robustness, and reproducibility of the results.
Insight 1: Real-world tool-use tasks remain challenging for current LMM agents. Despite multimodal capabilities, no model exceeds 50% in Goal Accuracy ($\mathbf{G_{acc}}$). o4-mini, the best performer, achieves only 45%, while most open-source models are below 30%, highlighting the difficulty of tool use and final answer consistency in real-world settings.
@@
Phi-4-VL-Instruct 0.13 0.21 0.24 0.61 0.19 0.47 0.40 0.11 0.26 0.42
InternVL2.5-8B 0.45 0.31 0.47 0.68 0.47 0.52 0.60 0.28 0.55 0.58
Gemma-3-4B 0.26 0.30 0.78 0.61 0.54 0.38 0.54 0.27 0.67 0.60
InternVL3-8B 0.46 0.34 0.54 0.68 0.45 0.70 0.40 0.20 0.59 0.62
VideoLLaMA3-7B 0.45 0.28 0.46 0.65 0.46 0.62 0.54 0.28 0.54 0.54
Qwen2.5-VL-7B 0.54 0.43 0.63 0.75 0.57 0.56 0.67 0.36 0.65 0.67
@@
Gemini-1.5-Pro 0.43 0.23 0.84 0.62 0.45 0.53 0.62 0.04 0.56 0.48
Gemini-2.5-Pro 0.40 0.36 0.81 0.72 0.48 0.64 0.73 0.40 0.56 0.62
GPT-4o 0.60 0.47 0.72 0.81 0.57 0.79 0.59 0.37 0.70 0.68
OpenAI-o4-mini 0.42 0.32 0.89 0.71 0.51 0.60 0.80 0.45 0.67 0.63
*Results for the additional models are in Appendix §D and cross-judge consistency study in Appendix §E.
Insight 2: Strong reasoning capabilities contribute to higher task success rates. Models that maintain consistently strong scores across reasoning metrics are likelier to perform well on final task outcomes. GPT-4o, for instance, achieves high scores in $\mathbf{F_{acc}}=0.81$, $\mathbf{F_{p}}=0.79$, and $\mathbf{S_{acc}}=0.59$, which correlates with a relatively high $\mathbf{G_{acc}}=0.37$ and $\mathbf{T^{s}_{acc}}=0.68$. Likewise, Qwen2.5, with solid reasoning scores ($\mathbf{C_{s}}=0.57$, $\mathbf{S_{acc}}=0.67$), attains $\mathbf{G_{acc}}=0.36$, outperforming many open-source peers. These results indicate that deep reasoning and structured execution enhance task success in vision-centric agents.
@@ 4.3 Error Analysis
Figure 4: Qualitative comparison of GPT-4o and VideoLLaMA3-7B on Agent-X visual reasoning tasks. GPT-4o hallucinates tool use and gives incorrect justifications; VideoLLaMA3-7B lacks temporal reasoning and frame alignment. More comparisons in Appendix §F.2.
Qualitative Analysis: We conduct a qualitative error analysis (Figure 4) to uncover key limitations in current vision-centric agentic models. Our findings reveal four prevalent issues: (1) agents often exhibit shallow reasoning by skipping frame-wise or multi-step analysis, especially in video-based tasks; (2) models frequently misuse or hallucinate tools not defined in the metadata, indicating poor tool grounding; (3) many responses violate expected output formats, producing incomplete or non-JSON-compliant traces that hinder evaluation; and (4) agents tend to hallucinate reasoning steps or bypass visual verification, resulting in factual inaccuracies. These insights highlight the need for better temporal reasoning, tool schema adherence, structured output enforcement, and visually grounded planning.
Table 5: Error Breakdown Across Models. Common planning, formatting, and reasoning errors on Agent-X across GPT-4o, Gemini-1.5-Pro, and InternVL3-8B. Formatting errors are counted alongside planning and reasoning errors. Extended details in Appendix §F.1.
Error Type GPT-4o Gemini-1.5-Pro InternVL3-8B
@@
Quantitative Analysis: We examine failure modes across three models (GPT-4o, Gemini-1.5-Pro, and InternVL3-8B), summarized in Table 5. Formatting errors are tracked separately to focus on reasoning ability. GPT-4o demonstrates strong structural awareness with fewer formatting errors (13.2% multiple tool calls, 6.7% final format), but often hesitates to act: 17.6% no response and 18.5% visual misinterpretation, indicating it plays safe but hesitates. Gemini-1.5-Pro is aggressive but format-fragile, with 44.5% JSON errors, 10.3% final formatting issues, and 34.3% visual misinterpretations; only 0.2% no response. InternVL3-8B shows a balanced yet challenged profile: 33.8% JSON errors, 16.4% final format errors, and struggles in spatial (13.5%) and visual reasoning (14.1%), suggesting it suffers both structural and perceptual issues.
@@ 5 Conclusion
We present Agent-X, a comprehensive benchmark for evaluating the reasoning capabilities of vision-centric agents. The tasks in Agent-X are context-rich and span diverse multimodal scenarios. Our evaluation framework includes executable tools categorized across six agentic environments, ensuring a wide range of task coverage. We evaluate more than 12 state-of-the-art LMMs on Agent-X and find that even strong closed-source models struggle with agentic tasks. Our insights provide actionable guidance for improving agentic capabilities. We believe that Agent-X will drive future research in enhancing multimodal reasoning and robust tool integration for vision-centric agents.
Limitations. While Agent-X covers six diverse environments, it is currently monolingual and may inherit certain distributional bias. The semi-automated approach to query and tool-chain generation improves efficiency but can occasionally produce lower-quality samples. Additionally, although Agent-X offers broad coverage, there remains ample opportunity for scalability.
@@
Section A Datacard for Agent-X
@@
Section D Evaluation of Additional Models
@@
Section H Instruction for Annotators
Section I Agent-X Task Examples
Section J Generation Prompt Details
Section K Evaluation Prompt Details
Section L Benchmark Curation Examples